At 2.30pm, David left Rose town for Lily town at a speed of 10km/h. At the same time, Muthu left Lily town for Rose town at a speed of 12km/h along the same road. If the distance between Rose town and Lily town was 6km, how far apart were the boys 15 minutes later?
This is what I know from the question.
David
Speed - 10km/h
Time - 36 minutes
Distance - 6km
Muthu
Speed - 12km/h
Time - 30 minutes
Distance - 6km

Answer by ramkikk66(644)


Question:
At 2.30pm, David left Rose town for Lily town at a speed of 10km/h. At the same time, Muthu left Lily town for Rose town at a speed of 12km/h along the same road. If the distance between Rose town and Lily town was 6km, how far apart were the boys 15 minutes later? 

Answer:
You have worked out how long it would take for each boy to reach the other town 
i.e. to cover the distance of 6 km. But the question only asks for how far they
are apart after 15 min.

David travels 10 km/hr.
So in 15 min he travels = 10*15/60 = 2.5 km
Muthu travels 12 km/hr
So in 15 min he travels = 12*15/60 = 3 km

Initially the boys are 6 km apart. Now since they are travelling towards each
other, the distance between them has reduced by (2.5 + 3) = 5.5 km.

So at the end of 15 min, the distance between them = 6 - 5.5 = 0.5 km.

Hope you got it. Let me know if you need further explanations :)